Apple0291:

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta: (2 points) 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture?

AngeI:

Hello welcome to QuestionCove c: So if A= pounds of Arabica beans then A=1,000 So you would do 700(1,000)+1,200R=1,000,000 You would then multiply the 700(1000) getting 700,000+1,200R=1,000,000 Next step would be to subtract 700,000 from both sides 700,000-700,000+1,200R=1,000,000-700,000 Which leads to 1,200R=300,000 You would then divide both sides by 1,200 to find out what R is R=250 So 250 pounds of Robusta beans should be in the mixture c: